# Structure and Dynamics of Magneto-Inertial, Differentially Rotating Laboratory Plasmas

### v1

@@ 1 Introduction
Although the MRI is commonly observed in both sheared-box and global disc numerical simulations, experimental demonstration has proven elusive. Existing laboratory experiments use the Taylor-Couette geometry to realise steady-state quasi-Keplerian, highly resistive MHD flows and study instabilities, turbulence and angular momentum transport in the nonlinear regime which develops gradually over hundreds of rotation periods. These experiments control the rotation profile from the edges of the flow, either by spinning the vessel containing a liquid, such as water or aqueous-glycerol (see Ji et al., 2006), or sodium or gallium alloys under an external axial or helical magnetic field (see Ji et al., 2001; Goodman & Ji, 2002; Hollerbach & Rüdiger, 2005; Stefani et al., 2006; Liu et al., 2006); or by applying electrical currents from the edge of Hall plasmas confined by permanent magnets at the boundary (Collins et al., 2012; Flanagan et al., 2020; Milhone et al., 2021). Recently, the Princeton MRI experiment have reported the detection of the standard MRI, i.e. with a purely vertical field, using a GaInSn eustetic alloy in a Taulor-Couette apparatus (Wang et al., 2022b, a).
@@
High-energy-density (HED) plasmas driven at pulsed-power and laser facilities offer access to a complementary regime, different from previous experiments. These plasmas are typically hot (electron temperature $T_{e}>20$ eV for pulsed-power and $T_{e}>100$ eV in laser-plasma experiments, respectively) and dense, with electron number densities $n_{e}>10^{18}$ cm-3 on both cases (see reviews by, e.g. Remington et al. (2006); Lebedev et al. (2019); Takabe & Kuramitsu (2021)). HED plasmas are strongly driven, exhibiting flow velocities $\mathbf{u}$ up to several hundreds of kilometers per second. As a consequence, their fluid and magnetic Reynolds numbers are much greater than unity and therefore are solidly in the ideal MHD regime Ryutov et al. (1999). Although HED facilities were originally designed for the inertial confinement fusion program, the last decades have seen the growth of a diverse ecosystem of laboratory astrophysics experiments looking at plasma jets, supernovae explosions, collisionless shocks, magnetic reconnection, fluctuation dynamo, and others (see Remington et al. (2006); Lebedev et al. (2019); Takabe & Kuramitsu (2021)). In addition, radiative cooling is often important both in this regime as well as in astrophysical accretion discs, and therefore HED laboratory experiments may open an avenue for exploring the effects of radiative cooling in a rotating plasma in conditions relevant to accretion discs which cannot be studied in liquid metal experiments.
@@
Nevertheless, the study of rotating plasmas in the HED regime remains a largely unexplored frontier. A handful of experiments investigating rotating plasmas have been fielded on pulsed-power generators (e.g. (Ampleford et al., 2008; Bennett et al., 2015)) alongside with numerical simulations (e.g. (Bocchi et al., 2013a, b)). Cvejić et al. (2022) found self-generated rotation in an imploding gas-puff Z pinch with knife edge electrodes. More recently, Valenzuela-Villaseca et al. (2023a) reported the discovery of a quasi-Keplerian rotation curve in a dedicated laboratory astrophysics platform, which adds a crucial piece for accurate modelling of accretion disc physics. On the theory side, Beresnyak et al. (2023) found evidence of the MRI in imploding, differentially rotating plasma simulations relevant to HED experiments. Beyond physics purely of interest to accretion discs, it has been theoretically discovered that, under rotation, the compressibility and heat capacity of gases (Geyko & Fisch, 2013) and plasmas (Geyko & Fisch, 2017) can be modified.
@@
In this article, we present a detailed experimental characterization of the plasma structure and dynamics on the Rotating Plasma Experiment (RPX) platform, extending results earlier reported by Valenzuela-Villaseca et al. (2023a). In particular, in this paper we study the wire array Z pinch ablation dynamics, plasma radial growth, structure and evolution of an axial jet launched from a rotating plasma column. Moreover, we present an in-situ determination of the flow velocity field on interferometry measurements. In addition, we present analytical models relevant to our experimental results regarding the ablating magnetic field, and evolution of the plasma radius and rotation velocity. The paper is structured as follows: Section 2 describes the experimental designed and diagnostics suite, together with an analytical model to calculate the ablation direction, and therefore allow controlling the total amount of initial angular momentum. Section 3 presents all experimental results, in particular 3.1 shows that, other than the re-direction of ablation flows (which introduce angular momentum), the ablation properties are identical to standard wire array Z pinches (Swadling et al., 2013; Harvey-Thompson et al., 2012a, b). In section 4 we show that the plasma exhibits a self-similar evolution, which is describable solely by conservation laws. Since our experimental results are in agreement with both our self-similar solution coupled and the calculated magnetic fields, these models can be combined to design future experiments without the need for computationally expensive numerical simulations. Finally, we use our self-similar solution to suggest future avenues for investigation.
@@ 2.1 The Rotating Plasma Experiment (RPX) pulsed-power platform
The experiments were conducted on the Mega-Ampere Generator for Plasma Implosion Experiments (MAGPIE) pulsed-power generator which delivers $1.4$ MA peak electrical current with $240$ ns rise-time (Mitchell et al. (1996)). The load is a wire array Z pinch (Benjamin et al., 1981; Aivazov et al., 1987, 1988; Deeney et al., 1995; Yadlowsky et al., 1996; Lebedev et al., 2000, 2001) consisting of a $R=8$ mm radius cylindrical arrangement of eight equally spaced aluminium wires (40 $\mu$m diameter each) and $10$ mm long. The wire array is surrounded by eight 1 mm thick stainless steel return posts, offset by $\varphi_{0}=13^{\circ}$ and positioned on a $R_{P}=11$ mm radius cylinder, as shown in Fig. 1a. The return posts are used to introduce angular momentum in the experiment (Bocchi et al., 2013a; Valenzuela-Villaseca, 2022; Valenzuela-Villaseca et al., 2023a).
@@
Ablating wire array Z pinches go through distinct phases (see review by Lebedev et al. (2005)) that are relevant to the RPX. In the ablation phase, the Lorentz force transports the coronal plasma inwards and the wires remain still acting as a mass reservoir. When a significant fraction of the mass has been ablated (empirically, between $50\%$ and $80\%$), breakages along the wires appear, which quickly change the load’s impedance, ending the ablation phase. During the following implosion phase, the current distribution is suddenly distributed and the remaining mass in the wires implodes in a shell-like implosion, snow-ploughing through the ablated trailing mass. In the stagnation phase the load’s mass reaches the axis, quickly decelerating and releasing an X-ray burst (e.g., Spielman et al. (1998)). We find that RPX is able to drive rotation throughout the ablation and implosion phases. If there is a final X-ray burst from the final stagnation of the rotating plasma is currently unknown.
@@
Figure 1: Schematic of experimental setup. a) End-on view of the experimental components together with the plasma dynamics in the plane of rotation. b) Side-on view. Top anode discs connects all wires and all return posts. Current path is split symmetrically through the load, but only one path through a wire and a return post is presented.
@@ 2.2 Diagnostics
The plasma is probed by a suite of diagnostics which encompasses laser interferometry imaging, self-emission imaging, and Thomson Scattering that allow multiple measurements of relevant plasma parameters, such as electron density, temperature, and velocity. On each experiment, the imaging diagnostics provide two orthogonal lines of sight namely vertical (along the rotation axis, end-on, xy-plane) and lateral (perpendicular to rotation axis, side-on, xz-plane) simultaneously. Both the diagnostics probing direction and line of sight are shown schematically in Figure 1.
@@ 3.1 Ablation dynamics and formation of rotating plasmas
Figure 3: Sequence of end-on optical self-emission images obtained in a single experiment. Emission from the outer edges are likely produced by some residual ablation of the experimental hardware. The self-emission intensity along the ablation flows decreases due to the combination of radiative cooling effects and a decreasing density towards the axis Valenzuela-Villaseca et al. (2023a).
On the axis, a circular bright ring is formed at around $130$ ns, corresponding to the rotating plasma. This formation time can be used to estimate the ablation velocity $V_{ab}\sim 70$ km/s, consistent with direct Thomson scattering measurements by Harvey-Thompson et al. (2012a) for standard wire array Z pinches. The self-emission structure exhibits reduced intensity on-axis. This is consistent with a dense plasma shell surrounding a low-density core. This has been observed in previous experiments by Bennett et al. (2015) also and is produced by the centrifugal barrier associated with dynamically significant angular momentum (Valenzuela-Villaseca et al., 2023b). Unfortunately, this sequence of images does not allow to estimate the rotation velocity, since the smooth ring characterizing the plasma has no discernible features that co-rotate with the plasma which can be tracked with time. At $\sim 170$ ns, radially-oriented features emerge from the ring. These are secondary shocks produced by the oblique interaction of the ablation flows undergoing thermal expansion (Swadling et al., 2013).
@@
Figure 4: End-on interferometry results obtained from the 355 nm wavelength probe. a) Raw interferogram. b) Processed line-integrated electron density map.
@@ 3.2 Propagation and density structure of axial jets
Fig. 5a shows a sequence of extreme ultra-violet (XUV) images of the launching and propagation of the axial jet. A $1$ $\mu$m-thick Mylar filter was used to remove soft XUV emission and improve the contrast. In these images, dark regions correspond to high plasma XUV self-emission intensity, whereas the light-coloured regions correspond to non-emitting metallic hardware and vacuum regions. The z-axis is oriented vertically with the position z = 0 defined as the approximate location of the top disc’s upper surface. Therefore, the wire array plasma is restricted to the region z < 0. Vertical light-coloured strips which correspond to the shadows of the return posts. The results show that the posts remain static throughout the experiment, with very little visible ablation and no overall deformations. This robustness is key to sustain a steady state off-axis ablation.
@@
The axial velocity of the flow can be estimated by following the upper end of the jet (“the tip”) through the sequence of images for the two XUV cameras fielded in the experiment. The result is presented in Fig. 5b, with each XUV camera presented in different colours. The horizontal uncertainty is given by the 5 ns of time integration, whereas the vertical error bars estimate the uncertainty in the position of the tip of the jet due to the smooth extinction of the plasma self-emission. The results show the linear trend of the position of the jet with time, from which a vertical velocity $u_{z}=100\pm 10$ km/s can be estimated.
Figure 5: Side-on XUV plasma self-emission data. a--f) Sequence of pinhole images obtained in the same experiment. Features corresponding to hardware components are annotated in italics. Panel (d) corresponds to the image presented in (Valenzuela-Villaseca et al. (2023a)), reprinted with the author’s permission. e) Length of axial jet, measured using the full height at half maximum intensity, as a function of time obtained as measured from the MCPs (labelled A and B) located on two opposing lines of sight. Image at $t=130$ ns is not presented and length corresponding to $t=240$ ns was not used in linear fit.
@@ 3.3 Thomson Scattering measurements of plasma flow velocity and temperature profile
The absence of inwards mass diffusion is consistent with the axial jet’s density profile being supported by a centrifugal barrier, which implies that it is spinning (Valenzuela-Villaseca et al., 2023b). In the experiments, direct measurements of the rotation and radial flow velocity components were obtained using an optical Thomson Scattering (TS) diagnostic. The TS probe beam passes horizontally through the rotating jet at $z=5.5$ mm above the top disc’s upper surface and passing approximately through the plasma rotation axis. We extend the analysis reported by Valenzuela-Villaseca et al. (2023a). In particular, we measure from the orthogonal interferometry line of sight the impact parameter which is important to interpret the flow velocity components as azimuthal and radial.
@@
Figure 8: Horizontal flow velocity field on end-on interferometry imaging. a) Full end-on laser interferometry (532 nm wavelength). Inner cathode diameter (dashed circular line), outer edge of rotating plasma (magenta dashed circle), plasma symmetry axis (solid magenta circle), and TS collection volumes (small white circles) are presented. b) Close-up to the rotating plasma region. Flow velocity field obtained from TS Doppler are indicated using arrows to scale. c) Calculated radial velocity for the inferred impact parameter ($b=250$ $\mu$m, blue circles) and zero impact parameter (same as Fig. 7 for reference. d) Azimuthal velocity. e) Specific angular momentum. f) Angular frequency.
@@ 3.3.1 Effect of the probe impact parameter on the velocity measurements
The result of the correction on the radial and azimuthal flow velocity components is shown in Fig. 8c and d, respectively. The result was discussed only schematically by Valenzuela-Villaseca et al. (2023a) to show that the identification of a quasi-Keplerian rotation curve is not dependent on the impact parameter. However, here we present a direct measurement using a diagnostic with a line of sight orthogonal to the TS geometry. We find that if this correction is ignored, then the result indicates that there is a constant side-ways velocity, however after accounting for the impact parameter allows correctly interpreting that near the axis $u_{\parallel}$ corresponds to $u_{\theta}$ rather than $u_{r}$. Although the magnitude of the rotation velocity changes, this correction may be important to correctly determine rotation curves close to the axis.
4 Discussion: self-similar radial evolution of rotating plasmas
The sequence of optical self-emission images allow tracking the radial evolution of the rotating plasma through time. As presented above, the plasma has a broadband self-emission distribution corresponding to a bright shell with reduced intensity on axis. Our goal is to use the information encoded in the radial dynamics to investigate the role of the ablation flows in the confinement and angular momentum injection in the experiment. In this section we use the data to analytically show that the radial evolution and rotation velocity exhibit a self-similar evolution, meaning that the radius and rotation velocity at a given time $t_{1}$ can be scaled to a different time $t_{2}$. To quantify the plasma radial dynamics, we have defined an inner and an outer characteristic plasma radii at half the height of the intensity peak of the shell, as indicated in Fig. 9a.

### v2

@@ 1 Introduction
Although the MRI is commonly observed in both sheared-box and global disc numerical simulations, experimental demonstration has proven elusive. Existing laboratory experiments use the Taylor-Couette geometry to realise steady-state quasi-Keplerian, highly resistive MHD flows and study instabilities, turbulence and angular momentum transport in the nonlinear regime which develops gradually over hundreds of rotation periods. These experiments control the rotation profile from the edges of the flow, either by spinning the vessel containing a liquid, such as water or aqueous-glycerol (see Ji et al., 2006), or sodium or gallium alloys under an external axial or helical magnetic field (see Ji et al., 2001; Goodman & Ji, 2002; Hollerbach & Rüdiger, 2005; Stefani et al., 2006; Liu et al., 2006); or by applying electrical currents from the edge of Hall plasmas confined by permanent magnets at the boundary (Collins et al., 2012; Flanagan et al., 2020; Milhone et al., 2021). Recently, the Princeton MRI experiment have reported the detection of the standard MRI, i.e. with a purely vertical field, using a GaInSn eustetic alloy in a Taylor-Couette apparatus (Wang et al., 2022b, a).
@@
High-energy-density (HED) plasmas driven at pulsed-power and laser facilities offer access to a complementary regime, different from previous experiments. These plasmas are typically hot (electron temperature $T_{e}>20$ eV for pulsed-power and $T_{e}>100$ eV in laser-plasma experiments, respectively) and dense, with electron number densities $n_{e}>10^{18}$ cm-3 on both cases (see reviews by, e.g. Remington et al. (2006); Lebedev et al. (2019); Takabe & Kuramitsu (2021)). HED plasmas are strongly driven, exhibiting flow velocities $\mathbf{u}$ up to several hundreds of kilometers per second. As a consequence, their fluid and magnetic Reynolds numbers are much greater than unity and therefore are solidly in the ideal MHD regime Ryutov et al. (1999). Although HED facilities were originally designed for the inertial confinement fusion program, the last decades have seen the growth of a diverse ecosystem of laboratory astrophysics experiments looking at plasma jets, supernovae explosions, collisionless shocks, magnetic reconnection, fluctuation dynamo, and others (see reviews by Remington et al. (2006); Lebedev et al. (2019); Takabe & Kuramitsu (2021)). In addition, radiative cooling is often important both in this regime as well as in astrophysical accretion discs, and therefore HED laboratory experiments may open an avenue for exploring the effects of radiative cooling in a rotating plasma in conditions relevant to accretion discs which cannot be studied in liquid metal experiments.
@@
Nevertheless, the study of rotating plasmas in the HED regime remains a largely unexplored frontier. Ryutov (2011) proposed colliding multiple laser-driven plasma jets to create a rotating torus that can expel jets along the poles. This concept has inspired work on pulsed-power plasmas, including this paper. In fact, a handful of experiments investigating rotating plasmas have been fielded on pulsed-power generators (e.g. (Ampleford et al., 2008; Bennett et al., 2015)) alongside with numerical simulations (e.g. (Bocchi et al., 2013a, b)). Cvejić et al. (2022) found self-generated rotation in an imploding gas-puff Z pinch with knife edge electrodes. More recently, Valenzuela-Villaseca et al. (2023) reported the discovery of a quasi-Keplerian rotation curve in a dedicated laboratory astrophysics platform, which adds a crucial piece for accurate modelling of accretion disc physics. On the theory side, Beresnyak et al. (2023) found evidence of the MRI in imploding, differentially rotating plasma simulations relevant to HED experiments. Beyond physics purely of interest to accretion discs, it has been theoretically discovered that, under rotation, the compressibility and heat capacity of gases (Geyko & Fisch, 2013) and plasmas (Geyko & Fisch, 2017) can be modified.
@@
In this article, we present a detailed experimental characterization of the plasma structure and dynamics on the Rotating Plasma Experiment (RPX) platform, extending results earlier reported by Valenzuela-Villaseca et al. (2023). In particular, in this paper we study the wire array Z pinch ablation dynamics, plasma radial growth, structure and evolution of an axial jet launched from a rotating plasma column. Moreover, we present an in-situ determination of the flow velocity field on interferometry measurements. In addition, we present analytical models relevant to our experimental results regarding the ablating magnetic field, and evolution of the plasma radius and rotation velocity. The paper is structured as follows: Section 2 describes the experimental designed and diagnostics suite, together with an analytical model to calculate the ablation direction, and therefore allow controlling the total amount of initial angular momentum. Section 3 presents all experimental results, in particular 3.1 shows that, other than the re-direction of ablation flows (which introduce angular momentum), the ablation properties are identical to standard wire array Z pinches (Swadling et al., 2013; Harvey-Thompson et al., 2012a, b). In section 4 we show that the plasma exhibits a self-similar evolution, which is describable solely by conservation laws. Since our experimental results are in agreement with both our self-similar solution coupled and the calculated magnetic fields, these models can be combined to design future experiments without the need for computationally expensive numerical simulations. Finally, we use our self-similar solution to suggest future avenues for investigation.
@@ 2 Experimental set up
Figure 1: Schematic of experimental setup. a) MAGPIE’s electrical current trace. b) End-on view of the experimental components together with the plasma dynamics in the plane of rotation. c) Side-on view. Top anode discs connects all wires and all return posts. Current path is split symmetrically through the load, but only one path through a wire and a return post is presented.
@@ 2.1 The Rotating Plasma Experiment (RPX) pulsed-power platform
The experiments were conducted on the Mega-Ampere Generator for Plasma Implosion Experiments (MAGPIE) pulsed-power generator which delivers $1.4$ MA peak electrical current with $240$ ns rise-time (Mitchell et al., 1996), as shown in Figure 1a. The load is a wire array Z pinch (Benjamin et al., 1981; Aivazov et al., 1987, 1988; Deeney et al., 1995; Yadlowsky et al., 1996; Lebedev et al., 2000, 2001) consisting of a $R=8$ mm radius cylindrical arrangement of eight equally spaced aluminium wires (40 $\mu$m diameter each) and $10$ mm long. The wire array is surrounded by eight 1 mm thick stainless steel return posts, offset by $\varphi_{0}=13^{\circ}$ and positioned on a $R_{P}=11$ mm radius cylinder, as shown in Fig. 1a. The return posts are used to introduce angular momentum in the experiment (Bocchi et al., 2013a; Valenzuela-Villaseca, 2022; Valenzuela-Villaseca et al., 2023).
@@
Ablating wire array Z pinches go through distinct phases (see review by Lebedev et al. (2005)) that are relevant to the RPX. In the ablation phase, the Lorentz force transports the coronal plasma inwards and the wires remain still acting as a mass reservoir. When a significant fraction of the mass has been ablated (empirically, between $50\%$ and $80\%$), breakages along the wires appear, which quickly change the load’s impedance, ending the ablation phase. During the following implosion phase, the current is suddenly redistributed and the remaining mass in the wires implodes in a shell-like implosion, snow-ploughing through the ablated trailing mass. In the stagnation phase the load’s mass reaches the axis, quickly decelerating and releasing an X-ray burst (e.g., Spielman et al. (1998)). We find that RPX is able to drive rotation throughout the ablation and implosion phases. If there is a final X-ray burst from the final stagnation of the rotating plasma is currently unknown.
@@ 2.2 Diagnostics
The plasma is probed by a suite of diagnostics which encompasses laser interferometry imaging, self-emission imaging, and Thomson Scattering that allow multiple measurements of relevant plasma parameters, such as electron density, temperature, and velocity. On each experiment, the imaging diagnostics provide two orthogonal lines of sight namely vertical (along the rotation axis, end-on, xy-plane) and lateral (perpendicular to rotation axis, side-on, xz-plane) simultaneously. Both the diagnostics probing direction and line of sight are shown schematically in Figure 1b and c, respectively. All diagnostics are timed relative to current start, which defines $t=0$.
@@ 3.1 Ablation dynamics and formation of rotating plasmas
Figure 3: Sequence of end-on optical self-emission images obtained in a single experiment. Emission from the outer edges are likely produced by some residual ablation of the experimental hardware. The self-emission intensity along the ablation flows decreases due to the combination of radiative cooling effects and a decreasing density towards the axis Valenzuela-Villaseca et al. (2023).
On the axis, a circular bright ring is formed at around $130$ ns, corresponding to the rotating plasma. This formation time can be used to estimate the ablation velocity $V_{ab}\sim 70$ km/s, consistent with direct Thomson scattering measurements by Harvey-Thompson et al. (2012a) for standard wire array Z pinches. The self-emission structure exhibits reduced intensity on-axis. This is consistent with a dense plasma shell surrounding a low-density core. This has been observed in previous experiments by Bennett et al. (2015) also and is produced by the centrifugal barrier associated with dynamically significant angular momentum (Valenzuela-Villaseca et al., 2024). Unfortunately, this sequence of images does not allow to estimate the rotation velocity, since the smooth ring characterizing the plasma has no discernible features that co-rotate with the plasma which can be tracked with time. At $\sim 170$ ns, radially-oriented features emerge from the ring. These are secondary shocks produced by the oblique interaction of the ablation flows undergoing thermal expansion (Swadling et al., 2013).
Figure 4: End-on interferometry results obtained from the 355 nm wavelength probe. a) Raw interferogram. b) Processed line-integrated electron density map. The contour of two secondary shocks have been highlighted with orange dotted lines.
@@ 3.2 Propagation and density structure of axial jets
Panels a–f in Fig. 5 shows a sequence of extreme ultra-violet (XUV) images of the launching and propagation of the axial jet. A $1$ $\mu$m-thick Mylar filter was used to remove soft XUV emission and improve the contrast. In these images, dark regions correspond to high plasma XUV self-emission intensity, whereas the light-coloured regions correspond to non-emitting metallic hardware and vacuum regions. The z-axis is oriented vertically with the position z = 0 defined as the approximate location of the top disc’s upper surface. Therefore, the wire array plasma is restricted to the region z < 0. Vertical light-coloured strips which correspond to the shadows of the return posts. The results show that the posts remain static throughout the experiment, with very little visible ablation and no overall deformations. This robustness is key to sustain a steady state off-axis ablation.
@@
The axial velocity of the flow can be estimated by following the upper end of the jet, which is representative of its length, through the sequence of images for the two XUV cameras fielded in the experiment. The result is presented in Fig. 5b, with each XUV camera presented in different colours. The horizontal uncertainty is given by the 5 ns of time integration, whereas the vertical error bars estimate the uncertainty in the position of the tip of the jet due to the smooth extinction of the plasma self-emission. The results show the linear trend of the position of the jet with time, from which a vertical velocity $u_{z}=100\pm 10$ km/s can be estimated.
Figure 5: Side-on XUV plasma self-emission data. a--f) Sequence of pinhole images obtained in the same experiment. Features corresponding to hardware components are annotated in italics. Panel (d) corresponds to the image presented in (Valenzuela-Villaseca et al. (2023)), reprinted with the author’s permission. e) Length of axial jet, measured using the full height at half maximum intensity, as a function of time obtained as measured from the MCPs (labelled A and B) located on two opposing lines of sight. Image at $t=130$ ns is not presented and length corresponding to $t=240$ ns was not used in linear fit.
@@ 3.3 Thomson Scattering measurements of plasma flow velocity and temperature profile
The absence of inwards mass diffusion is consistent with the axial jet’s density profile being supported by a centrifugal barrier, which implies that it is spinning (Valenzuela-Villaseca et al., 2024). In the experiments, direct measurements of the rotation and radial flow velocity components were obtained using an optical Thomson Scattering (TS) diagnostic. The TS probe beam passes horizontally through the rotating jet at $z=5.5$ mm above the top disc’s upper surface and passing approximately through the plasma rotation axis. We extend the analysis reported by Valenzuela-Villaseca et al. (2023). In particular, we measure from the orthogonal interferometry line of sight the impact parameter which is important to interpret the flow velocity components as azimuthal and radial.
@@
Figure 8: Horizontal flow velocity field on end-on interferometry imaging. a) Full end-on laser interferometry (532 nm wavelength). Inner cathode diameter (dashed circular line), outer edge of rotating plasma (magenta dashed circle), plasma symmetry axis (solid magenta circle), and TS collection volumes (small white circles) are presented. b) Close-up to the rotating plasma region. Flow velocity field obtained from TS Doppler are indicated using arrows to scale. c) Calculated radial velocity for the inferred impact parameter ($b=250$ $\mu$m, blue circles) and zero impact parameter (same as Fig. 7 for reference. d) Azimuthal velocity.
@@ 3.3.1 Effect of the probe impact parameter on the velocity measurements
The result of the correction on the radial and azimuthal flow velocity components is shown in Fig. 8c and d, respectively. The result was discussed only schematically by Valenzuela-Villaseca et al. (2023) to show that the identification of a quasi-Keplerian rotation curve is not dependent on the impact parameter. However, here we present a direct measurement using a diagnostic with a line of sight orthogonal to the TS geometry. We find that if this correction is ignored, then the result indicates that there is a constant side-ways velocity, however after accounting for the impact parameter allows correctly interpreting that near the axis $u_{\parallel}$ corresponds to $u_{\theta}$ rather than $u_{r}$. Although the magnitude of the rotation velocity changes, this correction may be important to correctly determine rotation curves close to the axis.
4 Discussion: comparison to the self-similar evolution of a thin rotating shell
The sequence of optical self-emission images allow tracking the radial evolution of the rotating plasma through time. As presented above, the plasma has a broadband self-emission distribution corresponding to a bright shell with reduced intensity on axis. Our goal is to use the information encoded in the radial dynamics to investigate the role of the ablation flows in the confinement and angular momentum injection in the experiment. In this section we derive a self-similar solution (meaning that the radius and rotation velocity at a given time $t_{1}$ can be scaled to a different time $t_{2}$) for a thin plasma shell driven and confined by the inflowing ablation flows. It is assumed that the shell-width $\delta$ and mean-radius $r$ of the plasma are such that $\delta/r\ll 1$ over their evolution. We compare the data with the model and find that the edge of the plasma follows the self-similar trajectory of the thin shell, and additionally, the measured rotation velocity is consistent with this simple mathematical description. To quantify the plasma radial dynamics, we have defined an inner and an outer characteristic plasma radii at half the height of the intensity peak of the shell, as indicated in Fig. 9a.